# SparC: Sparse Representation and Construction for High-Resolution 3D Shapes Modeling

### v1

@@ 4.2 Comparation Results
Watertight remeshing results. We evaluate both our watertight remeshing (serving as VAE ground truth) on the ABO collins2022abo , Objaverse objaverse , and Wild datasets, using Chamfer distance (CD), Absolute Normal Consistency (ANC), and F1 score (F1) as metrics. As Table 1 shows, our SparseCubes consistently outperforms prior pipelines dora ; hunyuan2 ; craftsman (reported under “Dora-wt” dora for brevity) across all datasets and metrics. Remarkably, our wt-512 remeshed outputs even exceed the quality of the wt-1024 results produced by previous methods. Fig. 4 presents qualitative comparisons: our approach faithfully preserves critical components (e.g., the car wheel) and recovers fine geometric details (e.g., shelving frames).
VAE reconstruction results. We further assess our SparConv-VAE reconstruction against TRELLIS trellis , Craftsman craftsman , Dora dora , and XCubes xcubes in Table 4.1. Across the majority of datasets and metrics, our SparConv-VAE outperforms these prior methods. Qualitative results in Fig. 4.1 illustrate that our VAE faithfully reconstructs complex shapes with fine details (all columns), converts open surfaces into double-layered watertight meshes (columns 1, 4, and 6), and reveals unvisible hidden internal structures (column 6).
Figure 5: Qualitative comparison of single-image-to-3D generation. Under the same architecture and model size trellis , the generator trained with our SparConv-VAE yields more detailed reconstructions than TRELLIS trellis . Best viewed with zoom-in.
Generation results. We also valid the effectiveness of our SparConv-VAE for generation by fine-tuning the TRELLIS trellis pretrained model. Under the same diffusion architecture and model size (see Fig. 5), our approach synthesizes watertight 3D shapes with exceptional fidelity and rich detail—capturing, for example, the sharp ridges of pavilion eaves, the subtle facial features of human figures, and the intricate structural elements of robots.
@@ 4.3 Ablation Studies
Training cost. Thanks to our modality-consistent design, SparConv-VAE converges less than two days—about four times faster than previous methods, i.e., sparse voxel–based TRELLIS trellis and vecset-based approaches dora ; craftsman each require roughly seven days to train.
VAE with 2D rendering supervision. We also investigate the effect of incorporating 2D rendering losses into our VAE by using the mask, depth, and normal rendering objectives. We find that adding 2D rendering supervision yields negligible improvement for our SparConv-VAE. This observation concurs with Dora dora , where extra 2D rendering losses were likewise deemed unnecessary for a 3D-supervised VAEs. We attribute this to the fact that sufficiently dense 2D renderings encode essentially the same information as the underlying 3D geometry—each view being a projection of the same 3D shape.
@@ 5 Conclusion
We introduce SparC, a unified framework that tackles two longstanding bottlenecks in 3D generation pipelines: topology-preserving remeshing and modality-consistent latent encoding. At its heart, SparseCubes transforms raw, non-watertight meshes into fully watertight surfaces at high resolution—retaining fine details and small components. Building on this, SparConv-VAE, a sparse-convolutional variational autoencoder with a self-pruning decoder, directly compresses and reconstructs our sparse representation without resorting to heavyweight attention, achieving state-of-the-art reconstruction fidelity and faster convergence. When coupled with latent diffusion (e.g., TRELLIS), SparC elevates generation resolution for downstream 3D asset synthesis. Together, these contributions establish a robust, scalable foundation for high-fidelity 3D generation in both virtual (AR/VR, robotics simulation) and physical (3D printing) domains.
Limitations. While our SparseCubes remeshing algorithm excels at preserving fine geometry and exterior components, it shares several drawbacks common to prior methods. First, it does not retain any original texture information. Second, when applied to fully closed meshes with internal structures, hidden elements will be discarded during the remeshing process.

### v2

@@ 4.2 Comparation Results
Watertight remeshing results. We evaluate both our watertight remeshing (serving as VAE ground truth) on the ABO [6], Objaverse [8], and Wild datasets, using Chamfer distance (CD), Absolute Normal Consistency (ANC), and F1 score (F1) as metrics. As Table 1 shows, our Sparcubes consistently outperforms prior pipelines [2, 39, 15] (reported under “Dora-wt” [2] for brevity) across all datasets and metrics. Remarkably, our wt-512 remeshed outputs even exceed the quality of the wt-1024 results produced by previous methods. Fig. 4 presents qualitative comparisons: our approach faithfully preserves critical components (e.g., the car wheel) and recovers fine geometric details (e.g., shelving frames).
VAE reconstruction results. We further assess our Sparconv-VAE reconstruction against TRELLIS [32], Craftsman [15], Dora [2], and XCubes [23] in Table 4.1. Across the majority of datasets and metrics, our Sparconv-VAE outperforms these prior methods. Qualitative results in Fig. 4.1 illustrate that our VAE faithfully reconstructs complex shapes with fine details (all columns), converts open surfaces into double-layered watertight meshes (columns 1, 4, and 6), and reveals unvisible hidden internal structures (column 6).
Figure 5: Qualitative comparison of single-image-to-3D generation. Under the same architecture and model size [32], the generator trained with our Sparconv-VAE yields more detailed reconstructions than TRELLIS [32]. Best viewed with zoom-in.
Generation results. We also valid the effectiveness of our Sparconv-VAE for generation by fine-tuning the TRELLIS [32] pretrained model. Under the same diffusion architecture and model size (see Fig. 5), our approach synthesizes watertight 3D shapes with exceptional fidelity and rich detail—capturing, for example, the sharp ridges of pavilion eaves, the subtle facial features of human figures, and the intricate structural elements of robots.
@@ 4.3 Ablation Studies
Training cost. Thanks to our modality-consistent design, Sparconv-VAE converges less than two days—about four times faster than previous methods, i.e., sparse voxel–based TRELLIS [32] and vecset-based approaches [2, 15] each require roughly seven days to train.
VAE with 2D rendering supervision. We also investigate the effect of incorporating 2D rendering losses into our VAE by using the mask, depth, and normal rendering objectives. We find that adding 2D rendering supervision yields negligible improvement for our Sparconv-VAE. This observation concurs with Dora [2], where extra 2D rendering losses were likewise deemed unnecessary for a 3D-supervised VAEs. We attribute this to the fact that sufficiently dense 2D renderings encode essentially the same information as the underlying 3D geometry—each view being a projection of the same 3D shape.
@@ 5 Conclusion
We introduce Sparc3D, a unified framework that tackles two longstanding bottlenecks in 3D generation pipelines: topology-preserving remeshing and modality-consistent latent encoding. At its heart, Sparcubes transforms raw, non-watertight meshes into fully watertight surfaces at high resolution—retaining fine details and small components. Building on this, Sparconv-VAE, a sparse-convolutional variational autoencoder with a self-pruning decoder, directly compresses and reconstructs our sparse representation without resorting to heavyweight attention, achieving state-of-the-art reconstruction fidelity and faster convergence. When coupled with latent diffusion (e.g., TRELLIS), Sparc3D elevates generation resolution for downstream 3D asset synthesis. Together, these contributions establish a robust, scalable foundation for high-fidelity 3D generation in both virtual (AR/VR, robotics simulation) and physical (3D printing) domains.
Limitations. While our Sparcubes remeshing algorithm excels at preserving fine geometry and exterior components, it shares several drawbacks common to prior methods. First, it does not retain any original texture information. Second, when applied to fully closed meshes with internal structures, hidden elements will be discarded during the remeshing process.